# Multiplicative polynomial equations in infinitely many variables

### v1

@@ 5. Open problems
    For what subfields $E$ of the complex numbers do we have “approximate finite implies infinite exact” for all infinite sets of polynomial equations with all scalars in $E$?
@@
    The ”finitely many implies infinitely many” paradigm applies to linear equations and multiplicative polynomial equations with bounded or sequentially bounded norms (Theorems 34, and 5). For equations of this special type, countably many implies finitely many. But there may be other classes of equations for which this paradigm does not hold, but a stronger condition (”countably many implies uncountably many”) is true. The problem is to determine if such ”new” classes of equations exist.
@@
Let $(p,q)$ be a conjugate pair and let $M>0$. Let $\mathbf{a}_{i}=(a_{i,j})_{j=1}^{\infty}\in\ell^{p}$ for all $i\in\mathbf{N}$. The following are equivalent.
@@
    $\|\mathbf{x}\|_{q}\leq M\operatorname{\qquad\text{and}\qquad}\sum_{j=1}^{\infty}a_{i,j}x_{=}b_{i}\qquad\text{for all $i\in\mathbf{N}$.}$
F. Riesz proved the equivalence of (b) and (c) and Abian-Eslami proved (a) implies (c).
@@
First we prove that (a) implies (b). For all $h_{1},\ldots,h_{r}\in\mathbf{R}$,
@@
We have $\mathbf{x}_{r}=(x_{r,j})_{j=1}^{\infty}\in\ell^{q}$ and so, by Hólder’s inequality,
@@
Because this inequality is valid for all $N$ and because $\lim_{N\rightarrow\infty}g(N)=0$, it follows that (a) implies (b).
@@
In particular, $f(\mathbf{a}_{i})=b_{i}$. Because
@@
for all $\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\in W$, the linear functional $f$ has norm $\|f\|\leq M$. By the Hahn-Banach theorem, there is a bounded linear functional $F$ on $\ell^{p}$ such that $F(\mathbf{w})=f(\mathbf{w})$ for all $\mathbf{w}\in W$ and $\|F\|\leq M$.
@@
$b_{i}=f(\mathbf{a}_{i})=F(\mathbf{a}_{i})=(\mathbf{a}_{i},\mathbf{x})=\sum_{j=1}^{\infty}a_{i,j}x_{j}.$

### v2

@@ 4. Approximate finite implies exact infinite
5. An application to Dirichlet series
Here is an example of a multiplicative polynomial. Let $D$ be a positive integer and let $D<q\leq\infty$. For $d\in\{1,\ldots,D\}$, let ${s_{d}}\in\mathbf{C}$ with $\Re({s_{d}})>(q-d)/q$. Then
$\Re\left(\frac{qs_{d}}{q-d}\right)>1$
and so the Riemann zeta function
$\zeta\left(\frac{q{s_{d}}}{q-d}\right)=\sum_{j=1}^{\infty}\frac{1}{j^{q{s_{d}}/(q-d)}}=\sum_{j=1}^{\infty}\left(\frac{1}{j^{s_{d}}}\right)^{q/(q-d)}$
converges absolutely. Consider the sequence
$\mathbf{a}_{d}=\left(a_{d,j}\right)_{j=1}^{\infty}=\left(\frac{1}{j^{s_{d}}}\right)_{j=1}^{\infty}\in\ell^{q/(q-d)}$
for all $d\in\{1,\ldots,D\}$, where
$a_{d,j}=\frac{1}{j^{s_{d}}}$
and the polynomial
$(\mathbf{a}_{d},\mathbf{x}^{d})=\sum_{j=1}^{\infty}a_{d,j}x_{j}^{d}=\sum_{j=1}^{\infty}\frac{x_{j}^{d}}{j^{s_{d}}}.$
Let $k\in\{1,\ldots,D\}$. For all
we define the multiplicative coefficient
$a_{\Delta,J}$ $=a_{(d_{1},d_{2},\ldots,d_{k}),(j_{1},j_{2},\ldots,j_{k})}=a_{d_{1},j_{1}}a_{d_{2},j_{2}}\cdots a_{d_{k},j_{k}}$
$=\frac{1}{j_{1}^{s_{d_{1}}}\cdots j_{k}^{s_{d_{k}}}}$
. Then
We obtain the $q$-multiplicative polynomial of degree at most $D$
$P(\mathbf{x})$ $=P(\mathbf{x},s_{1},\ldots,s_{D})$
This series converges absolutely for all $\mathbf{x}=(x_{j})_{j=1}^{\infty}\in\ell^{q}$.
For $D=1$ and $q=2$ and for complex numbers $s$ with $\Re(s)>1/2$, the multiplicative polynomial is the Dirichlet series
$L(\mathbf{x})=\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s}}.$
This series converges absolutely for all $\mathbf{x}=(x_{n})_{n=1}^{\infty}\in\ell^{2}$. The corresponding linear equation in infinitely many variables is simply
$\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s}}=b.$
Applying Theorem 2, we obtain the following result, which asserts the existence of a Dirichlet series taking prescribed values at a given infinite set of points in the half-plane $\Re(s)>1/2$.
Theorem 6.
Let $I$ be an infinite set and let $(b_{i})_{i\in I}$ and $(s_{i})_{i\in I}$ be sequences of complex numbers with $\Re(s_{i})>1/2$ for all $i\in I$. Consider the linear equation in infinitely many variables
$L_{i}(\mathbf{x})=\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s_{i}}}=b_{i}.$
Let $M>0$. If, for every finite subset $S$ of I and every $\varepsilon>0$, the finite set of linear inequalities
$\left|\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s_{i}}}-b_{i}\right|\leq\varepsilon\qquad\text{for all $i\in S$}$
has a solution $\mathbf{x}_{S,\varepsilon}\in\ell^{2}$ with $\|\mathbf{x}_{S,\varepsilon}\|_{2}\leq M$, then the infinite set of linear equations
$\sum_{n=1}^{\infty}\frac{x_{n}}{n^{s_{i}}}=b_{i}\qquad\text{for all $i\in I$}$
has an exact solution $\mathbf{x}\in\ell^{2}$ with $\|\mathbf{x}\|_{2}\leq M$.
@@ 6. Open problems
    The “finitely many implies infinitely many” paradigm applies to linear equations and multiplicative polynomial equations with bounded or sequentially bounded norms (Theorems 34, and 5). For equations of this special type, countably many implies finitely many. But there may be other classes of equations for which this paradigm does not hold, but a stronger condition (“countably many implies uncountably many”) is true. The problem is to determine if such “new” classes of equations exist.
@@
Let $(p,q)$ be a conjugate pair and let $M>0$. Let $\mathbf{a}_{i}=(a_{i,j})_{j=1}^{\infty}\in\ell^{p}$ for all $i\in\mathbf{N}$ and let $(b_{i})_{i\in\mathbf{N}}$ be a sequence of numbers. The following are equivalent.
@@
    $\|\mathbf{x}\|_{q}\leq M\operatorname{\qquad\text{and}\qquad}\sum_{j=1}^{\infty}a_{i,j}x_{j}=b_{i}\qquad\text{for all $i\in\mathbf{N}$.}$
F. Riesz [5] proved the equivalence of (b) and (c) and Abian-Eslami [1] proved (a) implies (c).
@@
First we prove that (a) implies (b). For all $h_{1},\ldots,h_{r}$,
@@
We have $\mathbf{x}_{r}=(x_{r,j})_{j=1}^{\infty}\in\ell^{q}$ and so, by Hölder’s inequality,
@@
This inequality is valid for all $N$ and $\lim_{N\rightarrow\infty}g(N)=0$. Therefore,
Thus, (a) implies (b).
@@
In particular, $f_{d}(\mathbf{a}_{i})=b_{i}$. Because
@@
for all $\sum_{i=1}^{r}h_{i}\mathbf{a}_{i}\in W$, the linear functional $f$ has norm $\|f\|\leq M$. By the Hahn-Banach theorem, there is a bounded linear functional $F$ on $\ell^{p}$ such that $F(\mathbf{w})=f_{d}(\mathbf{w})$ for all $\mathbf{w}\in W$ and $\|F\|\leq M$.
@@
$b_{i}=f_{d}(\mathbf{a}_{i})=F(\mathbf{a}_{i})=(\mathbf{a}_{i},\mathbf{x})=\sum_{j=1}^{\infty}a_{i,j}x_{j}.$